# ReMindRAG: Low-Cost LLM-Guided Knowledge Graph Traversal for Efficient RAG

### v1

@@ 5.2.2 Evaluation on System Cost-Efficiency and Memorization
Setups. To evaluate the cost efficiency (along with its effectiveness) of ReMindRAG, we conducted multiple sets of comparative experiments under identical conditions. Here, given a dataset $A$ containing documents and user queries, we consider three setups: (1) Same Query: The model initially has already been evaluated on dataset $A$ and has memorized information from $A$ and is subsequently re-evaluated again on the same dataset. (2) Similar Query: The model is re-evaluated again on a dataset $A^{\prime}$ whose queries are semantically equivalent paraphrases of those in $A$ (cf. Section˜C.3 for implementation). (3) Different Query: The model is re-evaluated again on a dataset $A^{\prime\prime}$ whose queries are distinct from those in $A$ but share similar questions (cf. Section˜C.4 for implementation). Importantly, of all the datasets we tested, only the Multi-Hop QA dataset satisfies the criteria for the Different Query setup [^7] We use the ”hotpot_dev_distractor” version, which contains intentionally designed distractors.. Additionally, to evaluate the memorization capability of ReMindRAG comprehensively, we further consider multi-turn memorization. This involves initially evaluating ReMindRAG  on the same dataset $A$ multiple times (i.e., memorizing $A$ multiple times).
@@ 5.3.2 Memory Stability
Table 4: Dimensional Truncation Experiment. The horizontal axis represents the dimension truncation ratio; for example, 1/2 indicates retaining the first half of the dimensions. The vertical axis denotes the number of updates. The values in the table correspond to the answer accuracy, and all experiments were conducted with the same default parameters.
@@ D.2 Compute Resources
In our experiments, we employed GPT-4o-mini [34] and Deepseek-V3 [28] as the underlying large language models (LLMs). (All operations utilizing large language models (GPT-4o, GPT-4o-mini, Deepseek-V3) are completed by invoking APIs). Notably, the proposed experiments can be readily replicated on ordinary machines, ensuring broad accessibility and reproducibility for the research community. Experiments were executed on a dedicated research workstation configured with an AMD Ryzen 7 7800X3D 8-Core Processor, an NVIDIA GeForce RTX 4070 Ti SUPER GPU, and $64$ GB of DDR5 RAM.
@@ E.1 Additional Analysis on ReMindRAG
Here, $|E_{1}\cap E_{2}|$ denotes the cardinality of the intersection (i.e., the count of edges shared by both search paths), and $|E_{1}\cup E_{2}|$ denotes the cardinality of the union (i.e., the total count of distinct edges present in either search path). A $S$ value closer to $1$ indicates a higher overlap between the two search paths (implying more similar underlying logic), while a value closer to 0 indicates greater divergence.
@@ Appendix F Theoretical Analysis
Assuming $\forall q_{1},q_{2}\in Q$, their angle satisfies $\angle(q_{1},q_{2})<\theta$, the corresponding point set $N$ satisfies $\forall p_{1},p_{2}\in N,d_{\mathbb{S}^{d-1}}(p_{1},p_{2})<\theta$, where $d_{\mathbb{S}^{d-1}}$ denotes the geodesic distance on $\mathbb{S}^{d-1}$.
According to Jung’s Theorem on Spheres [19, 7] in $\mathbb{S}^{d-1}$, there exists a ball $P=\mathbb{B}^{d-1}(c_{p},r_{p})$ containing $N$ whose geodesic radius $r_{P}$ satisfies:
@@
From Jung’s Theorem, the center $c_{p}$ of the covering ball $P$ lies within the convex hull of $N$. Therefore, the unit vector $c=\varphi^{-1}(c_{p})$ satisfies:

### v2

@@ 5.2.2 Evaluation on System Cost-Efficiency and Memorization
Setups. To evaluate the cost efficiency (along with its effectiveness) of ReMindRAG, we conducted multiple sets of comparative experiments under identical conditions. Here, given a dataset AA containing documents and user queries, we consider three setups: (1) Same Query: The model initially has already been evaluated on dataset AA and has memorized information from AA and is subsequently re-evaluated again on the same dataset. (2) Similar Query: The model is re-evaluated again on a dataset $A^{\prime}$ whose queries are semantically equivalent paraphrases of those in AA (cf. Section˜C.3 for implementation). (3) Different Query: The model is re-evaluated again on a dataset $A^{\prime\prime}$ whose queries are distinct from those in AA but share similar questions (cf. Section˜C.4 for implementation). Importantly, of all the datasets we tested, only the Multi-Hop QA dataset satisfies the criteria for the Different Query setup [^7] We use the ”hotpot_dev_distractor” version, which contains intentionally designed distractors.. Additionally, to evaluate the memorization capability of ReMindRAG comprehensively, we further consider multi-turn memorization. This involves initially evaluating ReMindRAG  on the same dataset AA multiple times (i.e., memorizing AA multiple times).
@@ 5.3.2 Memory Stability
Table 4: Dimensional Truncation Experiment. The horizontal axis represents the dimension truncation ratio. The vertical axis denotes the number of updates. The values in the table correspond to the answer accuracy, and all experiments were conducted with the same default parameters.
@@ D.2 Compute Resources
In our experiments, we employed GPT-4o-mini [34] and Deepseek-V3 [28] as the underlying large language models (LLMs). (All operations utilizing large language models (GPT-4o, GPT-4o-mini, Deepseek-V3) are completed by invoking APIs). Notably, the proposed experiments can be readily replicated on ordinary machines, ensuring broad accessibility and reproducibility for the research community. Experiments were executed on a dedicated research workstation configured with an AMD Ryzen 7 7800X3D 8-Core Processor, an NVIDIA GeForce RTX 4070 Ti SUPER GPU, and 6464 GB of DDR5 RAM.
@@ E.1 Additional Analysis on ReMindRAG
Here, $|E_{1}\cap E_{2}|$ denotes the cardinality of the intersection (i.e., the count of edges shared by both search paths), and $|E_{1}\cup E_{2}|$ denotes the cardinality of the union (i.e., the total count of distinct edges present in either search path). A $S$ value closer to 11 indicates a higher overlap between the two search paths (implying more similar underlying logic), while a value closer to 0 indicates greater divergence.
@@ Appendix F Theoretical Analysis
Assuming $\forall q_{1},q_{2}\in Q$, their angle satisfies $\angle(q_{1},q_{2})<\theta$, the corresponding point set NN satisfies $\forall p_{1},p_{2}\in N,d_{\mathbb{S}^{d-1}}(p_{1},p_{2})<\theta$, where $d_{\mathbb{S}^{d-1}}$ denotes the geodesic distance on $\mathbb{S}^{d-1}$.
According to Jung’s Theorem on Spheres [19, 7] in $\mathbb{S}^{d-1}$, there exists a ball $P=\mathbb{B}^{d-1}(c_{p},r_{p})$ containing NN whose geodesic radius $r_{P}$ satisfies:
@@
From Jung’s Theorem, the center $c_{p}$ of the covering ball $P$ lies within the convex hull of NN. Therefore, the unit vector $c=\varphi^{-1}(c_{p})$ satisfies: